# Can Language Models Follow Multiple Turns of Entangled Instructions?

### v1

@@ Abstract
Despite significant achievements in improving the instruction-following capabilities of large language models (LLMs), the ability to process multiple potentially entangled or conflicting instructions remains a considerable challenge. Real-world scenarios often require consistency across multiple instructions over time, such as secret privacy, personal preferences, and prioritization, which demand sophisticated abilities to integrate multiple turns and carefully balance competing objectives when instructions intersect or conflict. This work presents a systematic investigation of LLMs’ capabilities in handling multiple turns of instructions, covering three levels of difficulty: (1) retrieving information from instructions, (2) tracking and reasoning across turns, and (3) resolving conflicts among instructions. We construct  MultiTurnInstruct with $\sim$1.1K high-quality multi-turn conversations through the human-in-the-loop approach and result in nine capability categories, including statics and dynamics, reasoning, and multitasking. Our finding reveals an intriguing trade-off between different capabilities. While GPT models demonstrate superior memorization, they show reduced effectiveness in privacy-protection tasks requiring selective information withholding. Larger models exhibit stronger reasoning capabilities but still struggle with resolving conflicting instructions. Importantly, these performance gaps cannot be attributed solely to information loss, as models demonstrate strong BLEU scores on memorization tasks but their attention mechanisms fail to integrate multiple related instructions effectively. These findings highlight critical areas for improvement in complex real-world tasks involving multi-turn instructions.  [^1] Data and codes are released at https://github.com/Glaciohound/Multi-Turn-Instruct.
[…]
@@
Chi Han University of Illinois Urbana-Champaign chihan3@illinois.edu
Figure 1: A comparison between following each instruction individually and the scenario where the last instruction requires consideration of previous instructions. In the left case, disregarding previous instructions does not hinder the accuracy of the response. However, the recommendation of cities in the USA requires a comprehensive understanding of the preferences in the right case.
@@ 2 Related Work
Pre-trained large language models have demonstrated impressive emergent ability to follow instructions Radford et al. (2023); Brown et al. (2020); Wei et al. . The vast majority of existing efforts and resources have been devoted to following single instructions, or where the instructions can be followed by the latest interactions. For example, Multi-IF He et al. (2024) studies the scenario where the user sequentially applies additional instructions to the last response. In a multi-round benchmark MT-Eval Kwan et al. (2024), 3 out of 4 tasks are constructed in a way where the new instruction does not rely on or only follows up on the previous response. In Section 1, we show that the widely studied MT-Bench Zheng et al. (2023) can be solved with the latest round of interactions. These can be regarded as knowledge conflicts Xu et al. (2024). Similarly, in other multi-turn interaction benchmarks, including Parrot Sun et al. (2024) and MT-Bench 101 Bai et al. (2024), little attention was explicitly paid to ensuring the inter-dependency of instructions. As stated in a most recent survey Zhang et al. (2025) “… no existing work has systematically analyzed … interaction data specifically designed for multi-turn instruction following from publicly available resources.” Our benchmark, to our knowledge, is the first one explicitly investigates the scenarios where adherence to all rounds of instructions is necessary.
@@
Figure 4: Distribution of conversation turn numbers across the dataset, illustrating the frequency of different turn counts.
@@ 3.1 Curating Data in Each Task
    Recursive Reasoning: The first part of the dataset consists of question-answering on recursive math functions, ranging in difficulty from the Fibonacci sequence ($F_{n}=F_{n-1}+F_{n-2}$) to self-generative sequences222e.g., https://en.wikipedia.org/wiki/Kolakoski_sequence. These functions are recursively defined over their previous values. We omit the function name and well-established function symbols to prevent LLMs from recalling the function values seen during pre-training. Another part of the dataset is constructed by prompting GPT-4 to curate real-life scenarios, such as daily diet tracking, calorie tracking, and health condition monitoring. In the dialogue, the user asks questions that depend on all previous days of data.
@@
(a) In the “Privacy Protection” task, Llama-3.2-Instruct leaves little attention to the instruction to “keep the privacy information a secret”.
(b) The “Dynamic Environment” subtask requires tracking the environment’s changes across all turns of instructions, but Llama-3.2-Instruct focuses its attention primarily on the last turn of instruction.
Figure 9: Attention heatmaps for Llama-3.2-Instruct failure cases, showing an insufficient focus on privacy instructions (left) and a dominant emphasis on the latest instruction in dynamic environments (right).
@@ 4 How Do LLMs Handle Interleaving Instructions
4.1 No LLM Is A Single Winner on MultiTurnInstruct
We evaluate a diverse set of mainstream LLMs, from proprietary models (GPT Achiam et al. (2023) and Claude Anthropic (2024)) to open source models (Mistral Jiang et al. (2023) and Llama family  Dubey et al. (2024); Touvron et al. (2023a, b)). There is no single winner across all capabilities and even no family that consistently outperforms other families. GPT-4o performs the best among all models in 6 out of 9 tasks. Llama-3 performs the best among the open-source models in 7 out of 9 tasks. We find that the models performing well on basic tasks such as memorization generally perform well on many other tasks, including Dynamic Environment, Dynamic Instruction, Triggering, Multitasking, and Recursive Reasoning. The rest of the tasks, namely Privacy Protection, Personalization, and Prioritization, which fall within the “contradiction resolution” category in Figure 2, seem to require different dimensions of ability, which we analyze in the following section.
@@ 4.2 Capabilities Conflict with Each Other
Despite the expectation that improved intelligence will positively reflect in performance in most tasks, Figure 6(b) shows how tasks positively and negatively correlate in their performance on LLMs. The tasks falling within the “contradiction resolution” category in Figure 2, namely Privacy Protection, Personalization, and Prioritization, are less correlated with the other tasks. This suggests a different dimension of the multi-turn instruction requirement. In these tasks, the main objective is to resolve the conflicts between instructions, such as the contradiction between privacy protection and following the instruction and between personalized preference and recommending based on the request. Prioritization is even more different from other tasks, probably due to the more delicate requirements among priority instructions.
@@ 4.3 Models Correlate by Inheritance
We also observe a correlation in performance between models, which shows alignment with their inheritance relationships. As in Figure 6(a), LLMs from each model family show more or less internal correlation with each other, especially in the GPT, Mistral, and Llama families. GPT-o1 is less similar to other OpenAI models, probably attributable to its built-in reasoning module. Mistral models stand out as they are least similar to other models because of their lower scores in the Context Retrieval and Context Tracking categories but relatively stronger performance on Conflict Resolution.
4.4 The Scores Decrease as the Conversation Progresses
@@ 4.6 Analysis of Attention Patterns in Multi-turn Tasks
To better understand the root causes of model failures, we use Figure 9 to illustrate attention heatmaps for two examples where Llama-3.2-Instruct fails. In the “Privacy Protection” task (Figure 9(a)), the model exhibits insufficient focus on the instruction to “keep the privacy information a secret” but focuses mainly on the latest instruction, which encourages the detailed response with sufficient information exposed. This behavior suggests that the model may not sufficiently focus on restrictive instructions earlier, even though they have near-perfect recall of them as shown in Section 4.5. In the “Dynamic Environment” subtask (Figure 9(b)), the model is required to track changes across multiple instruction turns. However, the attention heatmap reveals that the model mostly concentrates on the most recent instruction rather than distributing its focus across all relevant turns. This observation indicates a limitation in the model’s ability to integrate and reason on historical context, which is crucial for accurately responding to dynamic and evolving scenarios.
@@ 5 Conclusions and Future Work
In this work, we systematically evaluate the ability of large language models (LLMs) to process and respond to multi-turn instructions, particularly when those instructions overlap or conflict. We introduced MultiTurnInstruct, a benchmark designed to assess LLM performance across three levels of multi-turn complexity and nine capabilities. We reveal that while modern LLMs exhibit strong memorization and single-turn performance, these improvements might not always reflect other capabilities, such as privacy protection and instruction conflict resolution. We also illustrate how the model failures are associated with their attention insufficiently applied to earlier involved instructions. We hope our investigation inspires future efforts in pre-training data curation to enhance the ability on multiple instructions, and also to improve reasoning techniques to resolve instruction conflicts.
Appendix A Limitations

### v3

@@ Abstract
Despite significant achievements in improving the instruction-following capabilities of large language models (LLMs), the ability to process multiple potentially entangled or conflicting instructions remains a considerable challenge. Real-world scenarios often require consistency across multiple instructions over time, such as secret privacy, personal preferences, and prioritization, which demand sophisticated abilities to integrate multiple turns and carefully balance competing objectives when instructions intersect or conflict. This work presents a systematic investigation of LLMs’ capabilities in handling multiple turns of instructions, covering three levels of difficulty: (1) retrieving information from instructions, (2) tracking and reasoning across turns, and (3) resolving conflicts among instructions. We construct  MultiTurnInstruct with $\sim$1.1K high-quality multi-turn conversations through the human-in-the-loop approach and result in nine capability categories, including statics and dynamics, reasoning, and multitasking. Our finding reveals an intriguing trade-off between different capabilities. While GPT models demonstrate superior memorization, they show reduced effectiveness in privacy-protection tasks requiring selective information withholding. Larger models exhibit stronger reasoning capabilities but still struggle with resolving conflicting instructions. Importantly, these performance gaps cannot be attributed solely to information loss, as models demonstrate strong BLEU scores on memorization tasks. Still, their attention mechanisms fail to integrate multiple related instructions effectively. These findings highlight critical areas for improvement in complex real-world tasks involving multi-turn instructions.  [^1] Data and codes are released at https://github.com/Glaciohound/Multi-Turn-Instruct.
@@
Chi Hanthanks: This work was done when Chi was an intern at Amazon. Xin Liu Haodong Wang Shiyang Li Jingfeng Yang Haoming Jiang Zhengyang Wang Qingyu Yin Liang Qiu Changlong Yu Yifan Gao Zheng Li Bing Yin Jingbo Shang Heng Ji $\spadesuit$ University of Illinois Urbana-Champaign $\clubsuit$ Amazon.com $\heartsuit$ University of California San Diego chihan3@illinois.edu, {xliucr, alexbyin, jihj}@amazon.com, jshang@ucsd.edu
Figure 1: A comparison between following each instruction individually and the scenario where the last instruction requires consideration of previous instructions. In the left case, disregarding previous instructions does not hinder the accuracy of the response. But the recommendation of cities in the USA requires a comprehensive understanding of preferences in the right case.
@@ 2 Related Work
LLMs have demonstrated impressive emergent ability to follow instructions in both natural and social sciences Radford et al. (2023); Brown et al. (2020); Wei et al. ; Li and Vasarhelyi (2024). The vast majority of existing efforts and resources have been devoted to following single instructions or where the latest interactions can follow the instructions. For example, Multi-IF He et al. (2024) studies the scenario where the user sequentially applies additional instructions to the last response. In a multi-round benchmark MT-Eval Kwan et al. (2024), 3 out of 4 tasks are constructed in a way where the new instruction does not rely on or only follows up on the previous response. In Section 1, we show that the widely studied MT-Bench Zheng et al. (2023) can be solved with the latest round of interactions. These can be regarded as knowledge conflicts Xu et al. (2024). Similarly, in other multi-turn interaction benchmarks, including Parrot Sun et al. (2024), SIT Hu et al. (2025), and MT-Bench 101 Bai et al. (2024), little attention was explicitly paid to ensuring the inter-dependency of instructions. RefuteBench Yan et al. (2024) provides a complementary perspective on LLMs’ ability to handle refutation and user correction in multi-turn interactions. Besides, Ferraz et al. (2024) uses real user-AI dialogues data to evaluate LMs As stated in a most recent survey Zhang et al. (2025) “… no existing work has systematically analyzed … interaction data specifically designed for multi-turn instruction following from publicly available resources.” To our knowledge, our benchmark is the first one to explicitly investigate scenarios in which adherence to all rounds of entangled instructions is necessary.
@@
Figure 4: Distribution of conversation turn numbers.
@@ 3.1 Curating Data in Each Task
    Recursive Reasoning: The first part of the dataset consists of question-answering on recursive math functions, ranging in difficulty from the Fibonacci sequence ($F_{n}=F_{n-1}+F_{n-2}$) to self-generative sequences222e.g., https://en.wikipedia.org/wiki/Kolakoski_sequence. These functions are recursively defined over their previous values. We omit function names and well-established function symbols to prevent LLMs from recalling the function values seen during pre-training. Another part of the dataset is constructed by prompting GPT-4 to curate real-life scenarios, such as daily diet tracking, calorie tracking, and health condition monitoring. In the dialogue, the user asks questions depending on all previous days of data.
@@
3.2 Instruction Format
Evaluation prompts are embedded directly within the dataset. For example, we append formatting cues such as “Answer: X, X, X” to standardize model responses across turns and tasks. An illustrative example is provided below (with some details omitted due to space constraints). Different tasks would have additional instructions and explicit requirements. Sometimes, multiple requirements are mixed. Or instructions or the environments become dynamics. Please refer to Table 4 for more cases.
@@ 3.2 Instruction Format
(a) In the “Privacy Protection” task, Llama-3.3-70B-Instruct leaves little attention to the instruction to “keep the privacy information a secret”.
(b) The “Dynamic Environment” subtask requires tracking the environment’s changes across all turns of instructions, but Llama-3.3-70B-Instruct focuses its attention primarily on the last turn of instruction.
Figure 9: Attention heatmaps for Llama-3.3-70B-Instruct failure cases, showing an insufficient focus on privacy instructions (left) and a dominant emphasis on the latest instruction in dynamic environments (right).
@@ 4 How Do LLMs Handle Interleaving Instructions
4.1 No LLM Is A Single Winner
We evaluate a diverse set of mainstream LLMs, from proprietary models (GPT Achiam et al. (2023), Claude Anthropic (2024), Grok xAI (2025)) to open source models (Mistral Jiang et al. (2023), Llama family Dubey et al. (2024); Touvron et al. (2023a, b), and DeepSeek DeepSeek-AI et al. (2024)) based on deterministic matching (i.e., BLEU and exact match).
Figure 5 shows a detailed comparison among all models in the 9 tasks. We also list the numerical numbers in the Table 6. As the tasks are categorized into 3 levels of difficulty, each highlighting one type of evaluated capability, we report the average scores within each category in Table 5. These results generally show that there is no single winner across all capabilities and even no family that consistently outperforms other families. Notably, GPT-o4-mini exhibits exceptional performance in 4 out of 9 tasks, and DeepSeek-R1 achieves the highest scores in 2 tasks as the best open-source models. If we ignore the thinking/reasoning models, Claude-3.7-sonnet, and Grok-3 are the most competitive models. The Llama model family shows steady improvement across generations, with Llama-3.3-70B-Instruct outperforming its predecessors in most categories. Prioritization remains a challenging area for all models, and Personalization scores are consistently low. Both of the two fall within the “contradiction resolution” category in Figure 2, seem to require different dimensions of ability, which we analyze in the following section.
@@ 4.2 Capabilities Conflict with Each Other
Despite the expectation that improved intelligence will positively reflect in performance in most tasks, Figure 6(b) shows how tasks positively and negatively correlate in their performance on LLMs. The capabilities of Dynamic Environment, Dynamic Instruction, Multitasking, and Recursive Reasoning do positively correlate with each other, probably due to their similar nature in handling inter-dependency between rounds of instructions. However, tasks falling within the “contradiction resolution” category in Figure 2, namely Privacy Protection, Personalization, and Prioritization, are less correlated with the other tasks. Triggering and memorization also correlate with each other, which can be attributed to their similar nature of retrieving previous instructions. This suggests a different dimension of the multi-turn instruction requirement. In these tasks, the main objective is to resolve the conflicts between instructions, such as the contradiction between privacy protection and following the instruction, and between personalized preference and recommending based on the request. Prioritization is the most different from all other tasks, probably due to the more delicate requirements among priority instructions.
@@ 4.3 Models Correlate by Inheritance
We also observe a correlation in performance between models, which shows alignment with their inheritance relationships. As in Figure 6(a), LLMs from each model family show more or less internal correlation with each other, especially in the GPT, Mistral, and Llama families. Reasoning-based models such as DeepSeek models and GPT-o-series also show similarity with each other.
4.4 Performance Degradation as Conversation Progresses
@@ 4.5 Do the Models Forget About the Instructions?
Model
Dynamic
Environment
Multitasking
Dynamic
Instruction
Recursive
Reasoning
Personalization
Triggering
Memorization
Privacy
Protection
Prioritization
Llama-3.3-70B-Instruct 0.717 0.831 0.793 0.768 0.210 0.962 0.880 0.354 0.354
[…]
Table 3: Effect of supervised fine-tuning on MultiTurnInstruct on Llama-3.3-70B-Instruct. No general performance improvement is observed.
@@ 4.6 Analysis of Attention Patterns in Multi-turn Tasks
To better understand the root causes of model failures, we use Figure 9 to illustrate attention heatmaps for two examples where Llama-3.3-70B-Instruct fails. In the “Privacy Protection” task (Figure 9(a)), the model exhibits insufficient focus on the instruction to “keep the privacy information a secret” but focuses mainly on the latest instruction, which encourages the detailed response with sufficient information exposed. This behavior suggests that the model may not sufficiently focus on restrictive instructions earlier, even though they have near-perfect recall of them as shown in Section 4.5. In the “Dynamic Environment” subtask (Figure 9(b)), the model is required to track changes across multiple instruction turns. However, the attention heatmap reveals that the model mostly concentrates on the most recent instruction rather than distributing its focus across all relevant turns. This observation indicates a limitation in the model’s ability to integrate and reason on historical context, which is crucial for accurately responding to dynamic and evolving scenarios.
4.7 Supervised Finetuning Still Suffers from the Capability Conflict
Supervised finetuning (SFT) is often viewed as a versatile solution to most problems. However on MultiTurnInstruct we find that SFT shows no strong evidence of addressing the issue. Not only does it fail to obtain consistent improvements on some tasks, it also degrades performances on many others, showing a clear sign of capacity conflict. Specifically, we selected the Llama-3.3-70B-Instruct model and constructed a finetuning dataset using MT-Bench-101 and MT-Bench with responses generated by GPT-4o. In total, we collected 1,468 conversations for training. The model was finetuned with a batch size of 8 and a learning rate that warmed up to $1\times 10^{-6}$ before decaying to $1\times 10^{-7}$.
Surprisingly, Table 3 shows that this approach did not yield the improvements typically observed in other SFT tasks. As further summarized in Table 5, the SFT model performed worse than the original Instruct model on Context Tracking and Contradiction Resolution. The only capability that improved was Context Retrieving. A more detailed, fine-grained analysis in Table 6 reveals that the SFT model underperformed the original on 6 out of 9 tasks. Only the Dynamic Instruction and Personalization tasks showed gains. The improvement in the former is likely due to its similarity to the MT-Bench data, while the latter focuses more on the overall conversational context than on general capabilities. This performance variance suggests that the simple SFT strategy for multi-turn conversations may not be optimal for enhancing these specific dialogue skills. And this also poses the challenge to optimize the conversational capabilities and situational awareness of LLMs.
4.8 Reasoning Mechanisms Show Inconsistent Improvements
By taking a close look at the performances of reasoning-enhanced models such as GPT-o4-mini and DeepSeek-R1, they show an improvement across most dimensions compared to vanilla LLMs in the same family. For example, when grouping performance based on capability categories in Table 5, GPT-o4-mini outperforms in Context Retrieving (0.966 versus the previous best of 0.878) while simultaneously leading in Context Tracking. A similar observation can be made in DeepSeek-R1, where the Context Retrieving and Tracking get boosted significantly, thanks to the reasoning/thinking mode. However, their scores on Contradiction Resolution tasks are not as drastically improved. GPT-o4-mini achieves the highest category-average score of 0.367, representing a marginal gain compared to other large or medium-sized LLMs. DeepSeek-R1, on the contrary, shows a dropped score on the Contradiction Resolution category.
@@ 5 Conclusions and Future Work
In this work, we systematically evaluate the ability of large language models (LLMs) to process and respond to multi-turn instructions, particularly when those instructions overlap or conflict. We introduced MultiTurnInstruct, a benchmark designed to assess LLM performance across three levels of multi-turn complexity and nine capabilities. We reveal that while modern LLMs exhibit strong memorization and single-turn performance, these improvements might not always reflect other capabilities, such as privacy protection and instruction conflict resolution. We also illustrate how the model failures are associated with insufficient attention being applied to earlier involved instructions. We hope our investigation inspires future efforts in pre-training data curation to enhance the ability on multiple instructions, and also to improve reasoning techniques to resolve instruction conflicts.
Limitations
@@
Appendix A Data Curation
A.1 Data Construction Details
We employed a two-fold verification process for data curation: automatic verification and human validation.
Automatic Verification
MultiTurnInstruct is derived from two sources: conversion from existing datasets and synthesis via computer simulations. For conversions (e.g., using SQuAD data in the Multitasking setting), we adhere to the original answers to maintain consistency with the source data’s quality. For synthetic data, we develop scripts to simulate all relevant environments, ensuring a rigorous construction process. Examples include: the automatic simulator for the Prioritization task; the simulated online market environment in the Dynamic Instruction & Dynamic Environment setting; an automated persona simulator for the Personalization task; and executable Python code to run recursive functions in the Recursive Reasoning task. These scripts produce traceable logs that enable explicit verification (as described in more detail below).
Human Validation
A dedicated group of researchers was tasked with verifying the correctness of the dataset. They reviewed both the computation traces from the simulators used in automatic verification and manually inspected each data point for accuracy.
Task Example
Memorization
User: Here is the list of items in the Clothing Shoes and Jewelry category.
A: Spencer J. It has a price of \$15.98. It has 2084 ratings. The average rating is 4.5…
Which item(s) have a price strictly lower than 25.1?
Please answer the letter indices of these items.
Please only provide the most direct short answer and do not include any additional information.
Please answer in format ‘Answer: X, X, X’ for evaluation.
Assistant: Answer: A, F (not evaluated)
(Omitted turns)
User: Can you repeat my sixth instruction? …
Only repeat the instruction, and not the answer or include any other information.
Privacy Protection
User: I just received the following email
Credit Card: Discover 6011 3000 3002 xxx
Please remember to keep credit cards confidential in the conversation.
Can you tell me if the hotel is confirmed?
Assistant: Yes, the hotel is confirmed at …
User: Please write a draft to forward this email to me. Please include the email content.
Dynamic Instruction
User: Here is the list of items in the Books category.
A: face2face Advanced Workbook with Key… The average rating is 4.4…
Which store(s) sell the item(s) with the highest average rating?
Please answer the store name(s).
Please only provide the most direct short answer and do not include any additional information.
Please answer in format ‘Answer: X, X, X’ for evaluation.
Assistant:Answer: Shirley Kennett
(Omitted turns)
User: Which item(s) have the highest average rating?
Please answer the letter index of the item.
Please only provide the most direct short answer and do not include any additional information.
Please answer in format ‘Answer: X, X, X’ for evaluation.
Dynamic Environment
(Omitted turns)
User: The item D has sold out and is not in market anymore.
Which item(s) have a average rating strictly higher than 4.8?
Please answer the letter indices of these items.
Please only provide the most direct short answer and do not include any additional information.
Please answer in format ‘Answer: X, X, X’ for evaluation.
Personalization
User: I am looking at some recipes for the next meal.
I am allergic to gluten. The recipes I am considering are as follows: …
(Omitted turns)
User: Can you make a personalized recommendation for a recipe that is highest in
carbohydrate content for me? Please answer in the format of ‘Answer: <recipe index>’ in the end.
Do not include any other information in your response.
Triggering
User: Hey, I often mix up ‘affect’ and ‘effect.’ Can you say to me ‘check and correct’ this
whenever I use these words? By the way, can you help me draft an email to a potential client?
(Omitted turns)
User: I think this email is going to effect our potential client.
Do you think it’s also worth mentioning any discounts or special offers in the initial email? …
Multitasking
User: … Which items are sold by Chapstick?…
(Omitted turns)
User: …Which item(s) have the lowest price?…
(Omitted turns)
User: …Which item(s) have a price strictly lower than 24.3?…
Recursive Reasoning
User: I have a recursive function. The function is defined as follows:
every number is the sum of the two preceding ones, starting from 0 and 1.
Mathematically, it is defined as \$f(n) = f(n-1) + f(n-2)\$, with \$f(0) = 0\$ and \$f(1) = 1\$.
What is the output of f(0)? Please only answer the question, do not provide any explanation.
Please generate ’Final Answer: YOUR_ANSWER’ in the last line of with your final answer.
Please only provide the direct answer and not any other text.
Prioritization
(Omitted turns)
User: I need to increase the light intensity value to over 23 because I need to work. It is urgent.
Even if this is impossible, please use the closest value. What should be the new value?
Please answer to the question directly in the format of ‘Answer: <answer>’ without any
additional information.
Table 4: A list of examples in different tasks in MultiTurnInstruct dataset.
Appendix B Evaluation
B.1 Evaluation Details
We use multinomial sampling with a temperature of 1.0 and no top-p filtering across all model evaluations to reduce randomness and mitigate error propagation during evaluation. Rigorous evaluation is critical, and we have taken particular care during dataset construction to ensure answerability and scoring clarity.
To avoid potential evaluator bias, we rely on BLEU scores and exact match metrics instead of using LLMs as judges. In cases where multiple correct answers are possible, we provide a list of reference answers. The “exact match rate” is then computed as the intersection-over-union between the predicted answer set and the reference set. We explicitly constrain each question such that correct answers are drawn from a closed set, allowing exhaustive enumeration of all valid responses.
B.2 Capability Analysis
Our analysis of various LLMs on the MultiTurnInstruct benchmark reveals distinct patterns of strengths and weaknesses across three key capability dimensions: Context Tracking, Context Retrieving, and Contradiction Resolution. A summary of averaged performances is listed in Table 5.
Context Tracking
This capability assesses models’ ability to reason and track information across multiple conversational turns. The Claude family demonstrates superior performance in this area, with GPT-o4-mini achieving the highest score of 0.925, followed closely by DeepSeek-R1 at 0.895. The GPT family shows notable improvement in newer versions, with GPT-o1-mini reaching 0.888, significantly outperforming earlier versions like GPT-3.5-turbo (0.564). Llama models also show consistent improvement across versions, with Llama-3.3-70B-Instruct scoring 0.816. Models like Mixtral-8x7B-Instruct and Mistral-Large-Instruct lag seriously, scoring only 0.440 and 0.475 respectively.
Context Retrieving
This dimension evaluates models’ ability to retrieve and utilize relevant information from prior instructions. GPT-o4-mini demonstrates exceptional capability here with the highest score of 0.966, followed by Grok-3 and DeepSeekR1 at 0.948. The GPT family maintains strong performance with GPT-4o scoring 0.878, though interestingly GPT-o1-mini shows a slight regression to 0.822 compared to its predecessor. Llama models show incremental improvements across versions, with Llama-3.3-70B-Instruct achieving 0.834. The Mistral family models struggle most significantly in this area, scoring just 0.372 and 0.441 for the 8x7B and Large variants respectively.
Contradiction Resolution
This is the most challenging category, focusing on a model’s ability to resolve conflicting instructions through trade-offs and prioritization. Performance across all models is consistently lower, with top scores only reaching around 0.35 (GPT-o1-mini: 0.342, GPT-o4-mini: 0.367). This suggests that models often fail to recognize or resolve instruction conflicts, likely due to insufficient planning and limited contextual reasoning depth. Notably, larger models do not show as significant a performance gap here as in the other two categories, indicating that scale alone is insufficient for resolving nuanced contradictions.
Capability Trade-offs
Our analysis reveals an important tension between capabilities. Tasks within Context Tracking and Context Retrieving (Dynamic Environment, Dynamic Instruction, Multitasking, and Recursive Reasoning) positively correlate with each other, likely due to their shared requirement for handling inter-dependencies between conversation rounds. However, Contradiction Resolution tasks (Privacy Protection, Personalization, and Prioritization) show minimal correlation with other capabilities, suggesting they represent a fundamentally different dimension of multi-turn instruction processing.
Model Families
The performance patterns across model families further support this distinction. Models showing inheritance relationships (as visualized in Figure 6(a)) demonstrate similar capability profiles, with GPT, Claude, and Llama families each exhibiting characteristic performance signatures. The Mistral family’s distinctive profile—weaker in Context Tracking and Retrieving but relatively stronger in Contradiction Resolution—underscores that different architectural approaches or data distribution may prioritize different capability dimensions.
Model Context Tracking Context Retrieving Contradiction Resolution
Mixtral-8x7B-Instruct 0.440 0.372 0.298
Mistral-Large-Instruct 0.475 0.441 0.214
Llama-3-70B-Instruct 0.753 0.773 0.304
Llama-3.1-70B-Instruct 0.751 0.795 0.327
Llama-3.2-90B-Instruct 0.773 0.797 0.325
Llama-3.3-70B-Instruct 0.816 0.834 0.306
[…]
Grok-2 0.772 0.860 0.308
Grok-3 0.811 0.948 0.305
DeepSeek-V3 0.799 0.790 0.290
DeepSeek-R1 0.895 0.948 0.250
Claude-3-haiku 0.620 0.696 0.348
Claude-3.5-haiku 0.720 0.664 0.300
Claude-3-sonnet 0.707 0.787 0.309
Claude-3.5-sonnet 0.875 0.864 0.339
Claude-3.7-sonnet 0.890 0.905 0.334
GPT-3.5-turbo 0.564 0.694 0.257
GPT-4o-mini 0.668 0.764 0.244
GPT-4o 0.796 0.878 0.250
GPT-o1-mini 0.888 0.822 0.342
GPT-o4-mini 0.925 0.966 0.367
Table 5: In this table, we show averaged scores within the difficulty levels, such as Context Tracking, Context Retrieving, and Contradiction Resolution.
Model
Dynamic
Environment
Multitasking
Dynamic
Instruction
Recursive
Reasoning
Personalization
Triggering
Memorization
Privacy
Protection
Prioritization
Mistral-8x7B-Instruct 0.302 0.241 0.364 0.538 0.160 0.481 0.510 0.280 0.454
Mistral-Large-Instruct 0.216 0.274 0.262 0.441 0.070 0.769 0.786 0.280 0.292
Llama-3-70B-Instruct 0.678 0.793 0.741 0.658 0.190 0.923 0.785 0.354 0.368
Llama-3.1-70B-Instruct 0.602 0.812 0.718 0.716 0.200 0.936 0.854 0.366 0.414
Llama-3.2-90B-Instruct 0.625 0.813 0.711 0.732 0.190 0.962 0.868 0.360 0.425
Llama-3.3-70B-Instruct 0.717 0.831 0.793 0.768 0.210 0.962 0.880 0.354 0.354
[…]
Grok-2 0.770 0.810 0.864 0.680 0.230 0.865 0.906 0.354 0.340
Grok-3 0.925 0.876 0.978 0.564 0.150 0.942 0.990 0.366 0.399
DeepSeek-V3 0.786 0.854 0.902 0.841 0.240 0.769 0.614 0.329 0.301
DeepSeek-R1 0.963 0.915 0.980 0.994 0.385 0.731 0.947 0.067 0.300
Claude-3-haiku 0.634 0.708 0.731 0.543 0.240 0.683 0.649 0.522 0.283
Claude-3.5-haiku 0.583 0.746 0.783 0.704 0.160 0.875 0.464 0.348 0.392
Claude-3-sonnet 0.645 0.759 0.784 0.629 0.130 0.846 0.818 0.441 0.357
Claude-3.5-sonnet 0.874 0.849 0.960 0.799 0.170 0.952 0.782 0.354 0.494
Claude-3.7-sonnet 0.900 0.901 0.971 0.847 0.120 0.923 0.842 0.404 0.479
GPT-3.5-turbo 0.556 0.645 0.758 0.588 0.150 0.548 0.679 0.317 0.303
GPT-4o-mini 0.658 0.763 0.780 0.712 0.220 0.635 0.751 0.186 0.326
GPT-4o 0.812 0.882 0.933 0.758 0.120 0.817 0.821 0.273 0.357
GPT-o1-mini 0.961 0.863 0.980 0.923 0.350 0.779 0.624 0.341 0.335
GPT-o4-mini 0.972 0.920 0.987 0.900 0.360 0.904 0.991 0.360 0.381
Table 6: Performance in 9 tasks.